# Principal stratification with $U$-statistics under principal ignorability

### v3

@@ 4.2 Multiply robust estimators
Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 3, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao (2017) and Theorem 3 of Jiang et al. (2022). As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies and is often built upon the orthogonal moment functions (Chernozhukov et al., 2018, 2022) or the EIFs. Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al., 2021). Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al., 2018). Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2023). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq N\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al. (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
@@
Consider a scenario where $n=10$ and $K=3$. We can first form the partition =$\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$. Then, the partition $\mathcal{P}$ is formed as
@@
For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are more straightforward as the Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
[…]
@@
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we can obtain the following cross-fitting estimators for the PGCEs:
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n,L}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n,L}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (15)
where
@@
$\mathbb{U}_{n,L}g_{h}^{s}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$
and
$\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
Remark 5
Estimators given in (15) are the type II estimator proposed in Chernozhukov et al. (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
$\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})=|P_{l}|^{-1}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\hat{\vartheta}_{k})=|I_{k}|^{-1}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}),$
on $P_{l}$ and $I_{k}$, respectively, where $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ are correspondingly estimated on $P_{l}^{c}$ and $I_{k}^{c}$. Then, define
$\overline{\mathbb{U}}_{L}g_{h}^{s}(\widehat{\theta})=L^{-1}\sum_{l=1}^{L}\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\overline{\mathbb{E}}_{K}\psi_{D_{z}}(\widehat{\vartheta})=K^{-1}\sum_{k=1}^{K}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\widehat{\vartheta}_{k}),$
and the estimator for the PCEs can be formed as
$\widehat{\tau}_{h}^{10}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{10}(\widehat{\theta})}{[\overline{\mathbb{E}}_{K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{00}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{11}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$
which are the type I debiased machine learning estimators. We focus on the type II estimator because it is generally recommended for its numerical stability in pooling intermediate estimates based on split data, especially in smaller samples.
@@
We now discuss the asymptotic properties of the debiased machine learning estimators in (15). Assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following result summarizes the efficiency properties of (15).
Theorem 4 (Efficiency under nonparametric working models)
@@
[…]
@@
[…]
@@
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 2.
@@
Theorem 4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al. (2022, Lemma 15) and Escanciano and Terschuur (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 4 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator, which has been commonly adopted in debiased machine learning literature (Vansteelandt and Dukes, 2022; Wang et al., 2024). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.

### v4

@@ 4.2 Multiply robust estimators
Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),\allowbreak p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, each estimator $\widehat{\tau}_{h,\mathrm{tr}}^{s}$, $s\in\{10,00,11\}$, is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. This triple robustness property of $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is obtained by separately investigating the properties of its numerator and the denominator. Specifically, the denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via a nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
We propose further improvement of the parametric multiply robust estimator through debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov et al., 2018) or EIFs, but allows efficient estimation of causal effects with data-adaptive nuisance estimators. Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates. Compared to parametric working models, leveraging machine learning-based working models for the nuisance functions is more advantageous in terms of robustness against potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting—is adopted. In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov et al., 2018). Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
We adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2022). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq n\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al. (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. We use the following example to demonstrate the definition of $\mathcal{I}$ and $\mathcal{P}$.
@@
Consider a scenario where $n=10$ and $K=3$. We can first form the partition $\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$. Then, the partition $\mathcal{P}$ is formed as
@@
For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are straightforward Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
$(I_{1},I_{2})=\{(1,4),(1,5),(1,6),(1,7),(2,4),(2,5),(2,6),(2,7),(3,4),\allowbreak(3,5),(3,6),(3,7)\}.$
@@
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we can obtain the following debiased machine learning estimator for $\tau_{h}^{10}$,
$\widehat{\tau}_{h,\mathrm{dml}}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},$ (14)
@@
where $\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$ and $\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\allowbreak\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k})$.
@@
The debiased machine learning estimators $\widehat{\tau}_{h,\mathrm{dml}}^{00}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ can be constructed in a similar fashion and are given in Section S5 of the Online Supplement. We now discuss the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$. We first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following Theorem 3 summarizes the convergence and efficiency properties of the debiased machine learning estimators.
Theorem 3 (Nonparametric efficiency)
@@
    $\widehat{\theta}_{l}$ is consistent for $\theta$ and $\widehat{\vartheta}_{k}$ is consistent for $\vartheta$, with $\mathbb{E}|g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})-g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|^{2}\xrightarrow{p}0$, and $\mathbb{E}|\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k})-\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})|^{2}\xrightarrow{p}0$, for all $l$ and $k$;
@@
    $\|\widehat{\theta}_{l}-\theta^{*}\|=o_{\mathbb{P}}(n^{-1/4})$ and $\|\widehat{\vartheta}_{k}-\vartheta^{*}\|=o_{\mathbb{P}}(n^{-1/4})$, for all $l$ and $k$, where $\|\cdot\|$ denotes the $\mathcal{L}_{2}$-norm;
@@
$\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
@@
Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al. (2022, Lemma 15) and Escanciano and Terschuur (2022, Lemma 4). Regulatory conditions in Theorem 3 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt and Dukes, 2022). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via nonparametric bootstrap.
4.4 Sensitivity analysis for monotonicity
To point identify PGCEs, a central assumption is monotonicity (Assumption 2) which rules out the $01$ stratum. This stratum may be relevant, for example, when comparing two active interventions with unknown survival benefits against non-mortality outcomes. We thus describe a sensitivity analysis strategy to evaluate the robustness of the proposed estimators in the event of departure from monotonicity. Following the approach in Ding and Lu (2016), we define the sensitivity function $\eta(X)$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(X)=\mathbb{P}(S=01|X)/\mathbb{P}(S=10|X)$, where $\eta(X)\in[0,\infty)$; the monotonicity holds if $\eta(X)=0$, but $\eta(X)>0$ implies the existence of the $01$ stratum. The PCGEs in this scenario are thus $\tau_{h,\eta}^{s}$ for $s\in\{10,01,00,11\}$. To proceed, we first notice that, for $\eta(X)\neq 1$, the principal scores can be obtained as:
$e_{10,\eta}(X)=\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{01,\eta}(X)=\frac{\eta(X)\{p_{1}(X)-p_{0}(X)\}}{1-\eta(X)},$ (15)
$e_{00,\eta}(X)=1-p_{0}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{11,\eta}(X)=p_{1}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},$
with $0\leq\eta(X)\leq 1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(X)\}$ for all $s$. When $\eta(X)=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$ and the principal scores are not identifiable (Jiang et al., 2022, Section S4.3). Thus, this boundary case was ruled out. The identification of the PGCEs without monotonicity also requires an alternative version of the principal ignorability (Assumption S1), which we provide in Section S7.1 of the Online Supplement.
With a fixed value of the sensitivity function $\eta(X)$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity. The EIFs and the explicit forms of the bias-corrected estimators are given in Section S7.1 of the Online Supplement. Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the following result (proofs provided in Section S7.2 of the Online Supplement).
Theorem 4 (Multiply robustness and efficiency of bias-corrected estimators)
Suppose that Assumptions 1 and S1 in Section S7.1 of the Online Supplement hold with a known $\eta(X)\neq 1$, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, for $s\in\{10,01,00,11\}$, the estimator $\widehat{\tau}_{h,\eta,\mathrm{tr}}^{s}$ is (a) triply robust such that it is consistent for $\tau_{h,\eta}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; and (b) asymptotically linear with influence function $\varphi_{h,\eta}^{s}$, thereby achieving the local semiparametric efficiency bound under $\mathcal{M}_{\pi+p+\mu}$.